A gun has fired at a moving target. The bullet has a muzzle velocity of 1200 ft/s. The gun and the target are both on level ground, with target moving 60 mph away from the gun have a distance of 30,000 ft from the gun the moment the bullet was fired. What is the angle of the projectile needed to hit the target?


1
Expert's answer
2021-08-23T04:48:55-0400

We are given "(x,y)= (30,000)" and; "v_o =1200ft \/s" .

We know

"x = v_ot cos \theta \\
y = v_ot sin \theta —16t^2"

From equation (1) (by putting values), we get 

"t=\frac{3000 } {1200 cos \theta}"

Using this value on second equation, we get;

"0 = t(v_o sin \theta —16t) \\
v_osin \theta -16t = 0 \\
sin 2 \theta = frac{1}{15}\\
\frac{1}{2}sin^{-1} (\frac{1}{15}) =1.91° \\
\implies 1.91° ≤ \theta ≤ (90 -1.91)°"